OpenStudy (anonymous):

Solve. 6x^4-11x^2+3=0.

OpenStudy (anonymous):

factor>> (3x^2 - 1 )(2x^2 - 3)=0

OpenStudy (anonymous):

3x^2=1 x^2=1/3 ------------------------------------- 2x^2=3 x^2=3/2